"A=\sqrt{S(S-a)(S-b)(S-c)}"

Make S the subject of the formula


1
Expert's answer
2021-07-05T14:42:16-0400

"A=\sqrt {S(S−a)(S−b)(S−c)}"

square both sides of the equation to get:

"A^2=(\sqrt {S(S−a)(S−b)(S−c)})^2"

"A^2={S(S\u2212a)(S\u2212b)(S\u2212c)}"

but "S={a+b+c\over 2}"

"\implies A^2={{a+b+c\over 2}({a+b+c\over 2}−a)({a+b+c\over 2}−b)({a+b+c\over 2}−c)}"


"={{a+b+c\over 2}({a+b+c-2a\over 2})({a+b+c-2b\over 2})({a+b+c-2c\over 2})}"


"={{a+b+c\over 2}({b+c-a\over 2})({a+c-b\over 2})({a+b-c\over 2})}"


Multiply both sides of the equation by "{2\over a+b+c}"


"\implies A^2 \times {2\over a+b+c}={2\over a+b+c}\times {{a+b+c\over 2}({b+c-a\over 2})({a+c-b\over 2})({a+b-c\over 2})}"


"\implies {2A^2\over a+b+c}={({b+c-a\over 2})({a+c-b\over 2})({a+b-c\over 2})}"


"\implies {2A^2\over a+b+c}={{(b+c-a)(a+c-b)(a+b-c)\over 8}}"


Cross multiply to get


"\implies 16A^2=(a+b+c)(b+c-a)(a+c-b)(a+b-c)"


make "(a+b+c)" subject of the formula to get


"(a+b+c)={16A^2\over (b+c-a)(a+c-b)(a+b-c)}"


Divide both sides by 2 to get


"{(a+b+c)\over 2}={16A^2\over 2 (b+c-a)(a+c-b)(a+b-c)}"


"\implies {(a+b+c)\over 2}={8A^2\over (b+c-a)(a+c-b)(a+b-c)}"


Recall that "S={a+b+c\over 2}"


"\therefore" "S={8A^2\over (b+c-a)(a+c-b)(a+b-c)}"